OpenStudy (anonymous):

f(x) =2(x + 4), find x if F(x)= 20

OpenStudy (anonymous):

Equal the equation to 20 2(x+4)=20

OpenStudy (anonymous):

x+4=10 x=16

OpenStudy (anonymous):

2(20+4)=48

OpenStudy (anonymous):

@doubledrive :It says find x when f(x)=20 ,Not Find F(20)... Please check up your answer and whoever who ticked as a best response....

OpenStudy (anonymous):

@romeoblue:My answer is the correct answer not doubledrive's .

OpenStudy (anonymous):

@Eyad x=6, not 16

OpenStudy (anonymous):

lol @thechocoluver445 :Yup! Sry its 6